If the marginal propensity to consume (MPC) is 0.75, and if the goal is to increase real GDP by $400 million, then by how much would government spending have to change to generate this increase in real GDP?

Group of answer choices

a. $200 million.

b. $400 million.

c. $140 million.

d. $100 million.
Answer #1

Ans) K = 1/(1-MPC) = 1/(1-0.75) = 4

We know, K =

= $400 million / 4 = $100 million.

Option d.

[ K is multiplier, Y is income and I is investment]